OpenStudy (anonymous):

A bag contains 6 red marbles, 4 blue marbles, 7 green marbles, and 3 yellow marbles. A marble is drawn from the bag and is not replaced. Then a second marble is drawn. What is the probability that both marbles drawn were red?

OpenStudy (anonymous):

6/20*5/19 = 30/380 = 3/38